OpenStudy (lena772):

Generate the first 5 terms of this sequence: f(1) = 1 and f(2) = 2, f(n) = f(n - 1) + f(n - 2), for n > 2. 1, 2, 3, 4, 5 1, 2, 3, 5, 7 1, 2, 3, 5, 8 1, 3, 5, 7, 9

OpenStudy (lena772):

Third one is the answer, right? @Easyaspi314 @jim_thompson5910

OpenStudy (lena772):

@Zale101

OpenStudy (anonymous):

f(3) = f(2) + f(1) = 2 + 1 = 3

OpenStudy (anonymous):

Continue the procedure to find f(4)...

OpenStudy (anonymous):

This is the Fibonacci sequence

OpenStudy (lena772):

f(4) = f(4 - 1) + f(4 - 2) f(4) = f(3) + f(2)

OpenStudy (lena772):

f(4) = 5

OpenStudy (lena772):

@Easyaspi314 Is the third one the answer? :/